# Multi-Turn Code Generation Through Single-Step Rewards

### v1

@@ 1 Introduction
    We propose a multi-turn Best-of-N (BoN) approach for inference-time scaling and present benefits of a learned verifier to select the code solution at each turn.
@@
    Our approach $\mu$Code outperforms leading multi-turn approaches on MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) benchmarks. Our ablations demonstrate that learned verifiers aid in learning better generators and show promising scaling law trends with higher inference budgets.
@@ 2 Background
We model this as a Markov Decision Process (MDP), where the state is the interaction history $s_{t}=\{x,y_{1},o_{1},\dots,y_{t-1},o_{t-1}\}$ where $s_{1}=\{x\}$, and the action is the code $a_{t}=y_{t}$. The oracle reward is defined as $R(s_{t},a_{t})=R(x,a_{t})=C(x,y_{t})$ if $y_{t}$ passes all public and private tests (terminating the episode), or 00 otherwise.
@@
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y_{t}\sim\pi_{\theta}(\cdot|s_{t})}\left[\sum_{t=1}^{T}\gamma^{t}R(x,y_{t})\right].$ (1)
@@ 3 $\mu$Code: Multi-turn Code Generation
5:     Construct a local search expert using verifier $\pi_{\star}^{i}(x)=\arg\max_{y\in\mathcal{D}(x)}R_{\phi}^{i}(x,y)$
@@ 3.2 Training Verifier
Binary Cross-Entropy loss (BCE): The nominal way to train the verifier is to directly predict the oracle reward  (Cobbe et al., 2021):
@@ 3.3 Training Generator
$\mu$Code comprises a generator $\pi_{\theta}$ trained to produce code solutions conditioned on the initial problem and execution observations from previous turns. Given a dataset $\mathcal{D}$, $\mu$Code iteratively trains the generator to find the optimal code solution labeled using the local expert over the learned verifier. For this step, $\mu$Code extracts all code solutions from $\mathcal{D}$ for every problem $x$. An expert is then created by picking the best solution, $y^{\star}$, which achieves the highest score using with the learned verifier $R_{\phi}(x,y)$ and is given by
$y^{\star}=\pi_{\star}(x)=\arg\max_{y\in\mathcal{D}(x)}R_{\phi}(x,y).$ (4)
Using this expert dataset, we relabel the dataset $\mathcal{D}$ with the optimal solutions for each prompt:
@@ 3.4 Inference: Multi-turn Best-of-N
The selected code $y_{t}^{*}$ is executed in the environment over public tests to obtain the execution feedback $o_{t}$. This solution and the feedback is provided as context to the generator at the next turn to repeat this procedure. The search ends once $y_{t}^{*}$ passes all public tests or when the turn limit is reached. Consequently, even if $R_{\phi}(\cdot)$ grants a high score to a code solution, inference continues until the solution has successfully cleared all public tests, thus mitigating potential errors by $R_{\phi}(\cdot)$. The final response $y^{*}_{t}$ is then passed through the oracle verifier to check its correctness. Algorithm 2 describes a description of multi-turn BoN. We found it beneficial to use the reward model trained with samples of the latest generator $\pi_{\theta}$ (see Table 1).
@@ 3.5 Analysis
A MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$ with horizon $T$ is one-step recoverable if the advantage function of the optimal policy $\pi^{*}$, defined as $A^{*}(s,a)=Q^{*}(s,a)-V^{*}(s)$, is uniformly bounded for all $(s,a)$, i.e. $A^{*}(s,a)\leq 1$.
@@ 4 Experiments
Through our experiments, we aim to analyze (1) How does $\mu$Code compare to other state-of-the-art methods? (2) Does the learned verifier help during training and inference-time? (3) Which loss function works better for learning a verifier?
@@ 4.1 Setup
We conduct experiments on MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) where the agent needs to generate code solutions in Python given natural language descriptions. We train the methods on the MBPP training set which comprises 374 problems and evaluate on the MBPP test set and HumanEval (HE) dataset which have 500 and 164 problems. We further describe the prompts and the split of public and private tests in Appendix A.3 and A.4.
@@
We compare $\mu$Code with single and multi-turn baselines. For the single and multi-turn settings, we report metrics by generating solutions from Llama models which we denote as Instruct. We also compare with STaR (Zelikman et al., 2022) where the correct solutions of the Instruct model are used for fine-tuning (FT). We also compare to a multi-turn version of STaR, called Multi-STaR. Here, we collect multi-turn rollouts using the Instruct model and use trajectories terminating in a correct code solution for FT. For multi-turn BoN search, we collect the solutions that pass public tests, and then we select the best one judged by a learned verifier. Note that this verifier is specifically trained for each generator.
@@
We measure the performance with the BoN accuracy, which quantifies the accuracy of the solution chosen by a verifier from $N$ candidate solutions. The generator is allowed $T=3$ turns and the final turn is used for evaluation over private tests. At each turn, the verifier ranks $N=5$ solutions (unless stated otherwise) provided by the generator. For the BoN performance, we sample with a temperature of $0.7$. We also report the accuracy of generating correct solutions via greedy decoding.
@@ 4.2 Results
In Table 1, we compare the proposed algorithm $\mu$Code with the baselines. Here, we first evaluate the generators using code generated via greedy sampling for each problem ($N=1$). This measures the accuracy of generating a correct solution with a turn limit of $T=3$. We observe that multi-turn methods (both Instruct and Multi-STaR) perform better than their single-turn variants showing the importance of incorporating execution feedback. Our approach $\mu$Code outperforms Multi-STaR across both benchmarks with 1B-sized model demonstrating the efficacy of training generators with data obtained with a learned verifier. To highlight, our method $\mu$Code with a 1B parameter model achieves 1.9% performance gains compared to Multi-STaR on the HumanEval dataset. With an 8B-sized model, $\mu$Code outperforms baselines on MBPP whereas there is a drop when compared on HumanEval.
Approach Llama-3.2-1B Llama-3.1-8B
N MBPP HE MBPP HE
@@
Instruct 1 38.9 29.3 58.9 60.4
    +BoN 5 48.5 34.3 68.1 61.2
Multi-STaR 1 36.7 33.5 57.7 59.8
    +BoN 5 47.9 39.6 68.6 63.2
$\mu$Code 1 37.9 35.4 62.3 57.9
    +BoN 5 50.7 41.7 68.8 62.2
Table 1: Comparison of our method $\mu$Code with baselines across MBPP and HumanEval datasets. $N=1$ denotes generating solutions with 0 temperature. The Best-of-N (BoN) accuracy is computed with $N=5$ candidate solutions at each where the public tests and learned verifier is used for selection. We observe that $\mu$Code outperforms competing methods based on Llama-3.2-1B-Instruct and Llama-3.1-8B-Instruct models. The best performance for each dataset and model-sized is highlighted in bold and similar performances (within 1%) are underlined.
We further evaluate the effect of having a verifier for BoN search during inference, where a learned verifier selects the most promising candidate solution at each turn. In Table 1, we observe that all algorithms can benefit with BoN search. Remarkably, $\mu$Code attains a performance gain of up to 12.8% with the multi-turn BoN approach compared to greedy. Moreover, $\mu$Code outperforms leading baselines with BoN search on MBPP and HumanEval datasets by 2.8% and 2.1% with 1B sized-model and performs comparably on 8B-sized model.
@@ 4.3 Analysis
To understand the improvements, we conduct a component-wise ablation study where we 1) check the effect of oracle and learned verifiers for relabeling to train the generator (4.3.1), 2) evaluate different generators trained with and without learned verifiers (4.3.2), 3) check which verifier performs better multi-turn BoN search at test-time (4.3.3), 4) assess scaling behaviors at inference time with number of candidate generations ($N$) at each turn (4.3.4), and 5) study the benefits of learned verifiers during evaluation (4.3.5).
4.3.1 Verifier for relabeling
We compare different verifiers for relabeling data for training the generator. In contrast to $\mu$Code where the learned verifier is used to relabel (LV), we compare with relabeling using a correct solution (attains an oracle reward $R=1$) for the corresponding prompt (OV). We also compare with a variant where the generator is fine-tuned over combinations of data relabeled with both the oracle verifier and the learned verifier (OV+LV). Here, we concatenate the FT dataset obtained using both LV and OV. In Figure 2, we present results with the 1B-sized models across benchmarks and observe that having corrections with the oracle verifier outcome does not perform well. However, relabeling with both verifiers OV+LV outperforms the OV variant, further demonstrating that gains in the generator learned by $\mu$Code are coming from relabeling with a learned verifier. Lastly, the LV variant performed best on MBPP and comparably on HumanEval dataset when compared with LV+OV.
Figure 2: Comparison of relabeling with learned verifier (LV) and oracle verifier (OV) with the 1B model. The variant OV+LV aggregates a dataset from both verifiers for fine-tuning the generator. Note that OV+LV performs better than OV. However, relabeling with LV outperforms on MBPP and performs comparably on HumanEval, thereby demonstrating the benefits of leveraging the learned verifier for training the generator.
Figure 3: Comparison of $\mu$Code and baselines with 1B models on the ability of the learned generator to incorporate execution feedback at each turn. We observe that $\mu$Code consistently improves the BoN accuracy across turns on both datasets, whereas the baselines show marginal improvements with turns.
4.3.2 Varying the generator
In this section, we compare the multi-turn agents where the generator is trained with an oracle verifier (Multi-STaR) or a learned verifier ($\mu$Code). We evaluate the ability of the trained generator to utilize execution feedback and improve the code response across turns. We report the BoN accuracy till a turn $t$, which denotes the BoN accuracy of obtaining a correct solution till turn $t$. In Figure 3, we present the results with 1B-sized models. We observe that BoN accuracy improves with turns for $\mu$Code whereas the baseline agents show marginal improvements with successive turns. We believe that using a learned verifier for relabeling improves the generator’s ability to generate solutions with high reward values, and indeed recover better at every turn by utilizing the execution feedback.
4.3.3 Verifier at test-time
@@
Random 34.4 23.0 59.3 57.9
LV 40.3 27.0 61.1 61.2
PT 48.6 31.9 67.2 60.4
PT+LV 48.5 34.3 68.1 61.2
Multi-STaR
Random 35.6 30.5 59.2 57.7
LV 39.8 31.9 61.2 62.8
PT 46.7 37.6 67.6 60.0
PT+LV 47.9 39.6 68.6 63.2
@@
Table 2: Comparing BoN with different ways of picking solutions at each turn for multi-turn BoN search using the 1B sized model. The hierarchical approach of using public test and learned verifier (PT+LV) outperforms only using either public tests (PT) or the learned verifier (LV). The best performance for each dataset and model-size is highlighted in bold and similar performances (within 1%) are underlined.
@@
In our experiments with multi-turn BoN (Table 2), we pick the best solution based on the outcome of public tests and the scores of the learned verifier. In this experiment, we study different verifiers for ranking the candidate solutions at each turn. We test with Random strategy where the policy randomly picks from the $N$ solutions at each step. We compare to the public tests (PT) outcome that picks any solution that passes the public test. Note that this involves evaluating all generated solutions at every turn with public tests. We also compare to selecting a solution based on scores obtained via the learned verifier only (LV). This is crucial as in certain applications such privileged information like public tests are not available and the agents can benefit from learned verifiers during inference. Lastly, we compare with the combination of public tests and leveraging the scores of the learned verifier to break ties at each turn (PT+LV).
@@
In Table 2, we compare Base, Multi-STaR and $\mu$Code with different verifiers at test-time. We observe that LV outperforms Random strategy which shows that a learned verifier indeed aids in selecting better solutions among generations. Given the benefits of learned verifiers for multi-turn BoN search, they can be a good alternative when public tests are not available. Furthermore, using the outcome of public tests (PT) performed better than using learned verifiers (LV) except on the HumanEval datset for 8B-sized model. We believe that this gap can be further reduced by learning more powerful verifiers and leave it for future research. Interestingly, the hierarchical approach (PT+LV) that uses the learned verifier to break ties on the outcomes of public tests performed best across methods and datasets. We hypothesize that using learned verifiers is beneficial in two scenarios. Firstly, if multiple solutions pass the public tests, then the learned verifier can filter out incorrect solutions which may not pass private tests. Secondly, if all candidate solutions are incorrect, then the learned verifier should choose the most promising solution at each turn. This is crucial as picking a better solution with the learned verifier can lead to more relevant feedback for recovering the true solution.
4.3.4 Test-time scaling

### v2

@@ 1 Introduction
    We propose a multi-turn Best-of-N (BoN) approach for inference-time scaling and present benefits of learned verifier to select the code solution at each turn.
@@
    Our approach $\mu$Code outperforms leading multi-turn approaches on MBPP (Austin et al., 2021), HumanEval (Chen et al., 2021) and CodeContests (Li et al., 2022a) benchmarks. Our ablations show that learned verifiers aid in learning better generators and show promising scaling law trends with higher inference budgets.
@@ 2 Background
We model this as a Markov Decision Process (MDP), where the state is the interaction history $s_{t}=\{x,y_{1},o_{1},\dots,y_{t-1},o_{t-1}\}$ where $s_{1}=\{x\}$, and the action is the code snippet $y_{t}$. The oracle reward is defined as $R(s_{t},y_{t})=R(x,y_{t})=C(x,y_{t})$ if $y_{t}$ passes all public and private tests (terminating the episode), or 00 otherwise.
@@
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y_{t}\sim\pi_{\theta}(\cdot|s_{t})}\left[\sum_{t=1}^{T}\gamma^{t}R(x,y_{t})\right],$ (1)
where $\gamma\in[0,1)$ is the discount factor. As shown in Eq. 1, the objective optimizes for a policy to generate the correct solution with as few turns as possible. However, at any step $t$, the agent can generate the correct code solution $y_{t}=y^{\star}$ such that $C(x,y^{\star})=1$ (as shown in Fig. 1 (a)) – a one-step recoverable process. In the following section, we describe $\mu$Code, a simple and scalable framework that leverages the one-step recoverability and reduces the problem of reinforcement learning to imitation learning.
@@ 3 $\mu$Code: Multi-turn Code Generation
5:     Construct a local search expert using verifier $\pi_{\star}^{i}(x)=\arg\max_{y\in\mathcal{D}(x)}\beta_{\textsf{O}}R(x,y)+\beta_{\textsf{L}}R_{\phi}(x,y)$
@@ 3.2 Training Verifier
Binary Cross-Entropy loss (BCE): The nominal way to train the verifier is to directly predict the oracle reward labels (Cobbe et al., 2021) as given by:
@@ 3.3 Training Generator
$\mu$Code comprises a generator $\pi_{\theta}$ trained to produce code solutions conditioned on the initial problem and execution observations from previous turns. Given a dataset $\mathcal{D}$, $\mu$Code iteratively trains the generator to find the optimal code solution labeled using the local expert over the learned verifier. For this step, $\mu$Code extracts all code solutions from $\mathcal{D}$ for every problem $x$. An expert is then created by picking the best solution, $y^{\star}$, which achieves the highest score using the learned verifier $R_{\phi}(x,y)$ when combined with the output of oracle verifier $R(x,y)$ and is given by
$y^{\star}=\pi_{\star}(x)=\arg\max_{y\in\mathcal{D}(x)}\beta_{\textsf{O}}R(x,y)+\beta_{\textsf{L}}R_{\phi}(x,y),$ (4)
where $\beta_{\textsf{O}}=1.0$ and $\beta_{\textsf{L}}=0.1$ denote the weights for oracle and learned rewards. Note that for creating the training dataset, we also use the ground labels from oracle verifier as they are available to the agent. The combination of both verifiers yielded better performance in our experiments. Using this expert dataset, we relabel $\mathcal{D}$ with the optimal solutions:
@@ 3.4 Inference: Multi-turn Best-of-N
The selected code $y_{t}^{*}$ is executed in the environment over public tests to obtain the execution feedback $o_{t}$. This solution and the feedback is provided as context to the generator at the next turn to repeat this procedure. The search ends once $y_{t}^{*}$ passes all public tests or when the turn limit is reached. Consequently, even if $R_{\phi}(\cdot)$ grants a high score to a code solution, inference continues until the solution has successfully cleared all public tests, thus mitigating potential errors by $R_{\phi}(\cdot)$. The final response $y^{*}_{t}$ is then passed through the oracle verifier to check its correctness. Algorithm 2 describes our propose procedure of multi-turn BoN search. We found it beneficial to use the reward model trained with samples of the latest generator $\pi_{\theta}$ (see Table 1).
@@ 3.5 Analysis
A MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$ with horizon $T$ is one-step recoverable if the advantage function of the optimal policy $\pi^{*}$, defined as $A^{*}(s,a)=Q^{*}(s,a)-V^{*}(s)$, is uniformly bounded for all $(s,a)$, i.e. $-1\leq A^{*}(s,a)\leq 0$.
@@ 4 Experiments
Through our experiments, we aim to analyze (1) How does $\mu$Code compare to leading state-of-the-art methods? (2) Does a learned verifier facilitate training a better generator? (3) Can the use of a learned verifier improve multi-turn BoN search at inference time? (4) Does the test-time search show scaling law trends? (5) Which loss function works better for learning a verifier for $\mu$Code?
@@ 4.1 Setup
We conduct experiments on MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) where the agent needs to generate code solutions in Python given natural language descriptions. We train the methods on the MBPP training set which comprises 374 problems and evaluate on the MBPP test set and HumanEval (HE) dataset which have 500 and 164 problems. We also compare methods on the DeepMind CodeContests dataset (CC, Li et al. (2022a)) where we train on 1000 problems sampled from the training set and evaluate on the 165 problems in the test set. We further describe the prompts and the split of public and private tests in Appendix C.1 and C.2. For training, we trained RFT and $\mu$Code for 2 iterations in MBPP and HumanEval datasets and for 1 iteration on CodeContests.
@@
We compare $\mu$Code with single and multi-turn baselines. For single and multi-turn settings, we report metrics with Llama-3.2-1B Instruct and Llama-3.1-8B-Instruct as base models. We also compare with rejection finetuning (RFT) where we collect multiple rollouts and filter trajectories with a correct solution for fine-tuning (FT). For multi-turn RFT, given a positive rollout $\{x,y_{1},o_{1},...,y_{T},o_{T}\}$, the model is finetuned over all the sub-trajectories $\{s_{i},y_{i+1}\}_{i=0}^{T-1}$. For the multi-turn BoN search at inference time, we used the verifier learned from the last iteration for $\mu$Code and trained a verifier with generated rollouts for the base and RFT models. Lastly, for multi-turn BoN search at inference time, we pick the best code solution with a hybrid approach where a solution passing public tests is preferred followed by ranking solutions with a learned verifier.
@@
We evaluate the methods by comparing the BoN accuracy. The generator is allowed upto $T=3$ turns and the final solution is used for evaluation over private tests. The BoN@1 evaluates the agent at producing the solution in the first attempt and is obtained via greedy decoding with a temperature of 00. The BoN accuracy measures the ability of verifiers to leverage test-time compute by generating $N$ candidate solutions in parallel at each turn. At each turn, the verifier ranks $N=5$ solutions (unless stated otherwise) provided by the generator. For the BoN performance, we sample with a temperature of $0.7$.
@@ 4.2 Results
In Table 1, we compare our proposed algorithm $\mu$Code with the baselines. We first evaluate the generators using code generated via greedy sampling for each problem (BoN with $N=1$). Our approach $\mu$Code outperforms RFT across both benchmarks with 1B-sized model demonstrating the efficacy of using one-step recoverability and learned verifiers. To highlight, our method $\mu$Code with the 1B model outperforms baselines by 2.2% and 1.3% on MBPP and HumanEval datasets. Interestingly, the performance of RFT drops when compared to the base Instruct model for the multi-turn setting which can be attributed to the fact that finetuning dataset consists of sub-trajectories with incorrect code solution at non-terminal steps. For the 8B-sized variant, we observe similar trends where we see that all algorithms benefit from the multi-turn BoN search at inference time. Additionally, $\mu$Code performs better than both single and multi turn baselines across benchmarks.
Method Llama-3.2-1B Llama-3.1-8B
N MBPP HE MBPP HE CC
@@
Instruct 1 35.1 31.1 60.3 59.7 4.8
    +BoN 5 47.3 35.7 69.7 62.9 13.8
RFT 1 31.1 31.7 58.9 61.2 7.2
    +BoN 5 46.7 34.1 68.4 62.8 14.9
$\mu$Code 1 37.9 35.4 62.1 60.9 7.9
    +BoN 5 51.1 41.5 70.6 63.8 16.3
Table 1: Comparison of our method $\mu$Code with baselines across MBPP, HumanEval, and CodeContests datasets. $N=1$ denotes generating solutions with 0 temperature. The Best-of-N (BoN) accuracy is computed with $N=5$ candidate solutions at each where the public tests and learned verifier are used for selection. We observe that $\mu$Code outperforms competing methods based on Llama-3.2-1B-Instruct and Llama-3.1-8B-Instruct models. The best performance for each dataset and model-sized is highlighted in bold and similar performances (within 1%) are underlined.
With more inference budget and multi-turn BoN search at test-time, where the learned verifier and outcomes of public tests are used to select the best solution at each turn, we observe that every method performs better with the test-time search procedure by upto 13%. For the 1B-sized models, our method $\mu$Code significantly outperforms baselines by 4.4% on MBPP and by 5.8% on Humaneval datasets. For the 8B-sized variant, $\mu$Code performs better than the baselines across datasets. On the more challenging task of CodeContests, all methods benefit from the BoN search and observe performance gains of upto 2$\times$. On this benchmark, $\mu$Code outperforms the RFT and base model baselines by 1.4%. We further investigate these trends with a Qwen model in Appendix B.1 and additional unit tests on MBPP and HumanEval benchmarks in Appendix B.2.
@@ 4.3 Analysis
To delve deeper into the improvements, we conduct a component-wise ablation study where we 1) check the effect of one-step recoverability and the learned verifier for creating the local expert (4.3.1), 2) compare different verifiers for multi-turn BoN search at test-time (4.3.2), 3) test the efficacy of agents at utilizing execution feedback (4.3.3), 4) assess scaling behaviors at inference time with number of candidate generations ($N$) at each turn (4.3.4), and 5) study different loss functions to train the verifiers (4.3.5).
4.3.1 What makes a good generator in $\mu$Code?
Method One-step Verifier MBPP HE
RFT Oracle 46.7 36.5
$\text{RFT}_{\textsf{LV}}$ Learned 49.0 38.9
$\mu\textsc{Code}_{\textsf{OV}}$ Oracle 48.2 38.0
$\mu\textsc{Code}_{\textsf{LV}}$ Learned 48.5 39.1
$\mu\textsc{Code}$ Both 51.1 41.5
Table 2: Comparison of using learned verifier (LV) and relabeling with one-step recoverability (One-Step) with the 1B-sized model. We observe that FT with the learned verifier $\text{RFT}_{\textsf{LV}}$ performs better than FT with the oracle verifier scores RFT. Moreover, $\mu$Code performs best when from both verifiers are used for relabeling.
Firstly, we compare different verifiers for training to demonstrate the benefits of using dense rewards obtained from learned verifiers. The RFT baseline uses the oracle verifier to filter trajectories and does not relabel subsequences for FT. In this experiment, we compare RFT to $\text{RFT}_{\textsf{LV}}$ where the learned verifier selects the top $K$ rollouts for each prompt ranked via the learned verifier for FT. We use $K=3$ in our experiments. We observe in Table 2 that finetuning with the learned verifier outperforms the RFT baseline. In contrast to RFT, which lacks training data for prompts with no correct solutions, $\text{RFT}_{\textsf{LV}}$ effectively trains the generator to ascend the dense rewards obtained from the learned verifier.
Secondly, we present the advantages of relabeling subtrajectories with our insight of one-step recoverability. We extend the baselines with our proposed relabeling strategy described in Sec. 3.3. We call these methods $\mu\textsc{Code}_{\textsf{OV}}$ ($\beta_{\textsf{O}}=1$, $\beta_{\textsf{L}}=0$) and $\textsc{Code}_{\textsf{LV}}$ ($\beta_{\textsf{O}}=0$, $\beta_{\textsf{L}}=1$) depending on the verifier used for relabeling. Table 2 shows that $\mu\textsc{Code}_{\textsf{OV}}$ outperforms the baseline RFT presenting the benefits of relabeling. The performance is similar for the setting where the learned verifier is used. To leverage the best of both worlds, $\mu$Code uses a linear combination of scores from learned and oracle verifier for relabeling which outperforms each variant by around 2%.
Figure 2: Compares the BoN performance at each turn (with 6 turns) on HumanEval, MBPP datasets. We also present results on a partially observable version of MBPP where we remove the public tests from the prompt to test the efficacy of methods at incorporating execution feedback– MBPP (POMDP). We observe that all methods improve performance with more turns on MBPP and HumanEval datasets. However, on MBPP (POMDP) the performance drops at first turn and $\mu$Code closes the gap with performance on MBPP compared to the baselines demonstrating its ability to incorporate execution feedback to improve code solutions at each turn.
4.3.2 Does Learned Verifier aid BoN Search?
@@
Random 31.7 24.6 58.0 57.9
LV 30.3 29.4 62.5 61.0
PT 46.4 33.4 68.4 60.7
PT+LV 47.3 35.7 69.7 62.9
RFT
Random 31.1 27.4 58.9 57.7
LV 33.2 29.6 62.2 61.3
PT 46.8 36.8 67.4 61.4
PT+LV 46.7 36.5 68.4 62.8
@@
Table 3: Comparing BoN with different ways of picking solutions at each turn for multi-turn BoN search using the 1B sized model. The hierarchical approach of using public test and learned verifier (PT+LV) outperforms picking solutions with only using either public tests (PT) or the learned verifier (LV). The best performance for each dataset and model-size is highlighted in bold and similar performances (within 1%) are underlined.
@@
We study the effect of different verifiers for ranking the candidate solutions to pick the best solution for multi-turn BoN search at inference time. We test with Random strategy where the policy randomly picks from the $N$ solutions. We compare to using the outcomes on public tests (PT) that picks any solution that passes the public test. Note that this involves evaluating all generated solutions at every turn with all the given public tests. We also compare to selecting a solution based on scores obtained via the learned verifier only (LV). This is crucial as in certain applications such privileged information like public tests are not available and the agents can benefit from learned verifiers to improve search during inference. Lastly, we compare with the combination of public tests and using the dense scores obtained from the learned verifier to break ties at each turn (PT+LV).
@@
In Table 3, we compare the baselines and $\mu$Code with different verifiers at inference-time. We observe that LV outperforms Random strategy which shows that a learned verifier indeed selects better solutions amongst the candidates. Furthermore, using the outcome of public tests (PT) performed better than using learned verifiers (LV) and performs similarly on the HumanEval datset for 8B-sized models. We believe that this gap can be further reduced by learning more powerful verifiers with larger datasets. Interestingly, the hierarchical approach (PT+LV) that uses the learned verifier to break ties on the outcomes of public tests performs best across methods and datasets. We hypothesize that using learned verifiers is beneficial in two scenarios. Firstly, if multiple solutions pass the public tests, then the learned verifier can filter out incorrect solutions which may not pass private tests. Secondly, if all candidate solutions are incorrect, then the learned verifier chooses the most promising solution at each turn. This is crucial as picking a better solution with the learned verifier can lead to more relevant feedback for recovering the true solution.
4.3.3 Can $\mu$Code utilize Execution Feedback?
We evaluate the ability of the trained generator to utilize execution feedback and improve the code response across turns. We report the BoN accuracy till a turn $t$, which denotes the accuracy of obtaining a correct solution within $t$ turns. Note that the agents were trained with rollouts of upto 3 turns and we report results with 6 turns for this experiment. In Fig. 2 (left and middle), we present the results with 1B-sized models where we observe that BoN accuracy improves with successive turns across the benchmarks.
To further understand if $\mu$Code learns to utilize execution feedback, we curated another test dataset from MBPP where we removed the sample unit tests provided in the prompt, and call it MBPP (POMDP). Removing public unit test information from the prompt makes the information from execution feedback essential and is similar to a partially observable MDP setting. In Fig. 2 (right), we compare the BoN accuracy over this evaluation dataset for $\mu$Code and the baselines. We observe that the performance at first turn drops for all methods by upto 13%. With the execution feedback at each turn, $\mu$Code improves the code solutions leading to gains of 15.9% from turn 1 to turn 6 and matches its performance on the MBPP dataset. In contrast, the base model and RFT were unable to close this gap in this partially observable setting and performed worse by around 6% that the MBPP dataset. This demonstrates the ability of $\mu$Code to recover better code solutions at each turn by utilizing the execution feedback at each turn, a trend not observed for the Instruct model and the RFT baselines.
4.3.4 Does $\mu$Code scale with inference budget?